Purchased 41,800 yards at a total cost of $434,720 Used 38,800 yards in producing 1,900 awnings Actual direct labor cost of $Standards: Direct materials 21.0 yards per awning at $11.00 per yard Direct labor 3.0 hours per awning at $12.00 per hour VarRequirement 1. Calculate the standard cost of one awning. Standard cost Standard cost per unit Direct materials Direct laborRequirement 2a. Calculate the direct material variances (Enter the variances as positive numbers. Enter currency amounts to tRequirement 2b. Calculate the direct labor variances. (Enter the variances as positive numbers. Enter currency amounts to theRequirement 2c. Calculate the variable manufacturing overhead variances. (Enter the variances as positive numbers. Enter currRequirement 2d. Calculate the fixed manufacturing overhead variances. (Enter the variance as a positive number. Label the varRequirement 3. Explain what each of the variances you calculated means and give at least one possible explanation for each ofVariable manufacturing overhead: Meaning Possible explanation Variance VOH rate VOH efficiency Fixed manufacturing overhead:Are any of the variances likely to be interrelated? The variance is likely to be related to the V variance. It is likely that
Answer #1

1) Standard Cost

Direct Materials $18*25yards =$450
Direct Labor $10* 5 hrs =$50
Variable OH $2* 5 hrs =$10
Fixed OH $9 * 5hrs =$45
STANDARD COST PER UNIT =$555

2)

(a) Direct Materials Variances

Standard Quantity(SQ) = 2400 awings * 25 Yards per unit = 60,000 Yards

Actual Quantity (AQ) = 59,100 Yards (Given)

Standard Price (SP) = $18

Actual Price (AP) = Total Cost/Total Yards Purchased = 1,104,432/63,840 = $17.3

Material Price Variance = (SP-AP)AQ = ($18-$17.3)59100 = 0.7 * 59100 = 41,370 F

Material Quantity Variance =(SQ-AQ)SP = (60000-59100)$18 = 900 * $18 = 16,200F

2)

(b)Direct Labor Variances

Standard Hours(SH) = 2400 awings * 5hrs per unit = 12,000hrs

Actual Hours (AH) = 11,860 hrs (Given)

Standard Rate (SR) = $10

Actual Rate(AR) = Total Cost/Total Hrs = 122,158/11,860 = $10.3

Direct Labor Variance = (SR-AR)AH = ($10-$10.3)11,860 = 0.3 * 11,860 = 3,358 A

Direct Labor Efficency Variance =  (SH-AH) SR = (12,000-11,860)$10 =140 * $10 =1400F

2)

(c)Variable Over Head Variances

Standard Rate(SR) = $2

Standard Hours(SH)= 2400 awings * 5hrs = 12,000hrs

Actual Hours (AH) =11,860hrs

Actual Rate (AR) = $27278/11860 =$2.3

Variable Over Head Rate Variance = ( SR-AR)AH = ($2-$2.3)11,860= 0.3*11,860 = 3,558 A

Variable Over Head Efficency Variance = (SH-AH) SR = (12,000-11,860) $2 = 140 * $2 =280F

,